Required information

[The following information applies to the questions displayed below.]

Kinnear Plastics manufactures various components for the aircraft and marine industry. Kinnear buys plastic from two vendors: Tappan Corporation and Hill Enterprises. Kinnear chooses the vendor based on price. Once the plastic is received, it is inspected to ensure that it is suitable for production. Plastic that is deemed unsuitable is disposed of.

The controller at Kinnear collected the following information on purchases for the past year:

Tappan Hill
Total purchases (tons) 5,500 8,500
Plastic discarded 165 425

The purchasing manager has just received bids on an order for 190 tons of plastic from both Tappan and Hill. Tappan bid $1,552 and Hill bid $1,539 per ton.

Required:

1-a. Assume that the average quality, measured by the amounts discarded from the two companies, will continue as in the past. What is the effective cost per ton for both Tappan and Hill? (Do not round intermediate calculations.)
Answer #1
Particulars Tappan Hill
Plastic required (tons) (A) 190 190
Bid per ton ($) (B) 1552 1539
Plastic discarded (% on past year purchases) (%) =(165/5500)*100 =(425/8500)*100
3% 5%
Plastic discarded (Tons) (C) = 190 * 3% = 190 * 5%
5.7 9.5
Total Cost (A*B) = D 294880 292410
Plastic excl. discarded (A - C) = E 184.3 180.5
Effective cost per ton D/E 1600 1620